OpenStudy (anonymous):

Find the value of k so that a line passing through (k, k – 3) and (–2, 4) has a slope of 1/4 A. –3 B. 2 C. 5 D. 10

OpenStudy (anonymous):

Can anyone please explain how to solve?

OpenStudy (luigi0210):

I got you!

OpenStudy (luigi0210):

You have to use the slope-intercept form formula: \[\frac{ 4-(k-3) }{ -2-k }=\frac{ 1 }{ 4 }\] Can you handle it from here?

OpenStudy (anonymous):

1-k / -2-k = 1/4

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Not exactly.. did you distribute your negative?

OpenStudy (anonymous):

7-k/-2-k = 1/4

OpenStudy (anonymous):

Is this correct?

OpenStudy (luigi0210):

There you go

OpenStudy (anonymous):

I don't know how to continue though:(

OpenStudy (luigi0210):

Cross multiply

OpenStudy (anonymous):

28-4k/ -2-k is what i got

OpenStudy (anonymous):

Is it correct?

OpenStudy (luigi0210):

good so far, keep going

OpenStudy (anonymous):

I am not sure should I divide or what?

OpenStudy (luigi0210):

\[\frac{ 7-k }{ -2-k }=\frac{ 1 }{ 4 }\] \[-2-k=28-4k\]

OpenStudy (anonymous):

does k=10?

OpenStudy (luigi0210):

Why yes, yes it does

OpenStudy (anonymous):

Thaaaank yooou:)